# Parameter Efficient Fine-tuning of Self-supervised ViTs without Catastrophic Forgetting

### v1

@@ 1 Introduction
Humans excel at continual learning, gradually acquiring new information while retaining previously learned concepts, with a complete loss of prior knowledge being a rare occurrence. In contrast, artificial neural networks, including vision image transformers (ViT) [8], often suffer from catastrophic forgetting of old concepts as new ones are learned [9, 18]. Catastrophic forgetting directly stems from the stability-plasticity dilemma, a fundamental issue in neural networks that balances the integration of new knowledge (plasticity) against the preservation of existing knowledge (stability) [19]. Addressing this dilemma could enhance their applicability in dynamic, real-world scenarios.
@@
PEFT strategies have been extensively explored in Natural Language Processing (NLP) [11, 26] and multimodal foundation models [24, 30, 25], garnering significant attention for their ability to fine-tune models without extensively altering their structure. The exploration of these strategies has recently extended to vision models [28], with research focusing on various aspects of parameter-efficient fine-tuning including efficiency [29, 13, 10], scalability [5, 17], transferability [22, 15], and robustness [4, 21]. On the other hand, efforts to tackle catastrophic forgetting have been made independently [31, 16, 23], yet incorporating cutting-edge PEFT techniques from NLP into the vision domain to mitigate catastrophic forgetting in Vision Transformers (ViTs) remains an underexplored area. Our research distinguishes itself from these broad explorations of PEFT characteristics by specifically focusing on leveraging PEFT techniques to mitigate catastrophic forgetting during the fine-tuning of Vision Transformers (ViTs), thereby presenting a novel study in this domain.

### v2

@@ 1 Introduction
Humans excel at continual learning, gradually acquiring new information while retaining previously learned concepts, with a complete loss of prior knowledge being a rare occurrence. In contrast, artificial neural networks, including vision image transformers (ViT) [8], often suffer from catastrophic forgetting of old concepts as new ones are learned [9, 19]. Catastrophic forgetting directly stems from the stability-plasticity dilemma, a fundamental issue in neural networks that balances the integration of new knowledge (plasticity) against the preservation of existing knowledge (stability) [20]. Addressing this dilemma could enhance their applicability in dynamic, real-world scenarios. In these cases, the model must continually be fine-tuned with data from new class or domain distributions  [27, 24, 18, 25], while preserving its prior knowledge.
@@
PEFT strategies have been extensively explored in Natural Language Processing (NLP) [11, 29] and multimodal foundation models [26, 33, 28], garnering significant attention for their ability to fine-tune models without extensively altering their structure. The exploration of these strategies has recently extended to vision models [31], with research focusing on various aspects of parameter-efficient fine-tuning including efficiency [32, 13, 10], scalability [5, 17], transferability [22, 15], and robustness [4]. On the other hand, efforts to tackle catastrophic forgetting have been made independently [34, 16, 23], yet incorporating cutting-edge PEFT techniques from NLP into the vision domain to mitigate catastrophic forgetting in Vision Transformers (ViTs) remains an underexplored area. Our research distinguishes itself from these broad explorations of PEFT characteristics by specifically focusing on leveraging PEFT techniques to mitigate catastrophic forgetting during the fine-tuning of Vision Transformers (ViTs), thereby presenting a novel study in this domain.